# Not All Language Model Features Are Linear

### v1

@@ Abstract
Recent work has proposed the linear representation hypothesis: that language models perform computation by manipulating one-dimensional representations of concepts (“features”) in activation space. In contrast, we explore whether some language model representations may be inherently multi-dimensional. We begin by developing a rigorous definition of irreducible multi-dimensional features based on whether they can be decomposed into either independent or non-co-occurring lower-dimensional features. Motivated by these definitions, we design a scalable method that uses sparse autoencoders to automatically find multi-dimensional features in GPT-2 and Mistral 7B. These auto-discovered features include strikingly interpretable examples, e.g. circular features representing days of the week and months of the year. We identify tasks where these exact circles are used to solve computational problems involving modular arithmetic in days of the week and months of the year. Finally, we provide evidence that these circular features are indeed the fundamental unit of computation in these tasks with intervention experiments on Mistral 7B and Llama 3 8B, and we find further circular representations by breaking down the hidden states for these tasks into interpretable components.
@@ 1 Introduction
Language models trained for next-token prediction on large text corpora have demonstrated remarkable capabilities, including coding, reasoning, and in-context learning [7, 1, 3, 45]. However, the specific algorithms models learn to achieve these capabilities remain largely a mystery to researchers; we do not understand how language models write poetry. Mechanistic interpretability is a field that seeks to address this gap by reverse-engineering trained models from the ground up into variables (features) and the programs (circuits) that process these variables [37].
One mechanistic interpretability research direction has focused on understanding toy models in detail. This work has found multi-dimensional representations of inputs such as lattices [30] and circles [27, 35], and has successfully reverse-engineered the algorithms that models use to manipulate these representations. A separate direction has identified one-dimensional representations of high level concepts and quantities in large language models [15, 29, 17, 6]. These findings have led to the linear representation hypothesis: that all representations in pretrained large language models are one-dimensional lines, and that we can understand model behavior as nonlinear manipulations of these linear representations [38, 6]. In this work, we bridge the gap between these two regimes by providing evidence that language models also use multi-dimensional representations.
@@ 1.1 Contributions
    In Section 3, we generalize the one-dimensional definition of a language model feature to multi-dimensional features, provide an updated superposition hypothesis to account for these new features, and analyze the reduction in a model’s representation space implied by using multi-dimensional features. We also develop a theoretically grounded and empirically practical test for irreducible features and run this test on some sample distributions.
@@
    In Section 4, we present a method for finding multi-dimensional features using sparse autoencoders and identify multi-dimensional representations automatically in GPT-2 and Mistral 7B, including circular representations for the day of the week and month of the year. To the best of our knowledge, we are the first to find an emergent circular representation in a large language model.
@@
    In Section 5, we propose two tasks, modular addition in days of the week and in months of the year, that we hypothesize will cause models to use these circular representations. We perform intervention experiments on Mistral 7B and Llama 3 8B to show that models do indeed use circular representations for these tasks. Finally, we present novel methods for decomposing LLM hidden states and reveal circles in the computed day of the week and month of the year.
@@ 2 Related Work
Linear Representations: Early word embedding methods such as GloVe and Word2vec, although only trained using co-occurrence data, contained directions in their vector spaces corresponding to semantic concepts, e.g. the well-known f(king) - f(man) + f(woman) = f(queen) [32, 40, 31]. Recent research has found similar evidence of linear representations in sequence models trained only on next token prediction, including Othello board positions [36, 26], the truth value of assertions [29], and numeric quantities such as longitude, latitude, birth year, and death year [15, 17]. These results have inspired the linear representation hypothesis [38, 11] defined above. Recent theoretical work provides evidence for this hypothesis, assuming a latent (binary) variable-based model of language [22]. Empirically, dictionary learning has shown success in breaking down a model’s feature space into a sparse over-complete basis of linear features using sparse autoencoders [6, 9]. These works assume that the number of linear features stored in superposition exceeds the model dimensionality [11].
Nonlinear Representations: There has been comparatively little research on nonlinear features. One recent paper [24] proves that a one-layer swapped order (MLP before attention) transformer can in-context-learn a nonlinear mapping function followed by a linear regression, implying that the “features” between the MLP and attention blocks are nonlinear. Another work [42] finds that a transformer trained on a hidden Markov model uses a fractal structure to represent the probability of each next token. These works analyze toy models and it is not clear if large language models will have similar nonlinear features. A separate idea [4] argues for interpreting neural networks through the polytopes they split the input space into, and identifies regions of low polytope density as “valid” regions for a potential linear representation. Finally, recent work on dictionary learning [6] has speculated about multi-dimensional feature manifolds; our work is most similar to this direction and develops the idea of feature manifolds theoretically and empirically .
@@
Circuits: Circuits research seeks to identify and understand circuits, subsets of a model (usually represented as a directed acyclic graph) that explain specific behaviors [37]. The base units that form a circuit can be layers, neurons [37], or sparse autoencoder features [28]. The first circuits-style work looked at the InceptionV1 image model and found line features that were combined into curve detection features [37]. More recent work has examined language models, for example the indirect object identification circuit in GPT-2 [47]. Given the difficulty of designing bespoke experiments, there has been increased research in automated circuit discovery methods [28, 8, 44].
@@
Interpretability for Arithmetic Problems: Prior work studies models trained on modular arithmetic problems $a+b=c\ (\mathrm{mod}\ m)$ and finds that models that generalize well have circular representations for $a$ and $b$ [27]. Further work shows that models use these circular representations to compute $c$ via a “clock” algorithm [35] and a separate “pizza” algorithm [49]. These papers are limited to the case of a small model trained only on modular arithmetic. Another direction has studied how large language models perform basic arithmetic, including a circuits level description of the greater-than operation in GPT-2 [16] and addition in GPT-J [43]. These works find that to perform a computation, models copy pertinent information to the token before the computed result and perform the computation in the subsequent MLP layers. Finally, recent work [14] investigates language models’ ability to increment numbers and finds linear features that fire on tokens equivalent modulo $10$.
3 Definitions and Theory
@@
In this section, we focus on $L$ layer transformer models $M$ that take in token input ${\bf t}=(t_{1},\ldots,t_{n})$, have hidden states $\mathbf{x}_{1,l},\ldots,\mathbf{x}_{n,l}$ for layers $l$, and output logit vectors $\mathbf{y}_{1},\ldots,\mathbf{y}_{n}$. Given a set of inputs $T$, we let $X_{i,l}$ be the set of all corresponding $\mathbf{x}_{i,l}$. This section focuses on hypotheses that describe how to decompose hidden states $\mathbf{x}_{i,l}$ into sums of functions of the input (features). While this is always possible if $M$ is deterministic via the “trivial” evaluation of $M$ itself, we are interested in decomposable, interpretable hypotheses for the construction of $\mathbf{x}_{i,l}$. We write matrices in capital bold, vectors and vector valued functions in lowercase bold, and sets in capital non-bold.
@@ 3.1 Multi-Dimensional Features
We define a $d_{f}$-dimensional feature of sparsity $s$ as a function ${\mathbf{f}}$ that maps a subset of the input space of probability $01-s>$ 0 into a $d_{f}$-dimensional point cloud in $\mathbb{R}^{d_{f}}$. We say that a feature is active on the aforementioned subset.
As an example, let $n=1$ (so inputs are single tokens) and consider a feature ${\mathbf{f}}$ that maps integer tokens to equispaced points in $\mathbb{R}^{1}$. Then ${\mathbf{f}}$ is a $1$-dimensional feature that is active on integer tokens, and if integer tokens occur $1\%$ of the time across the input distribution, ${\mathbf{f}}$ has sparsity $s=0.99$.
For features to be meaningful, they should be irreducible. Here, we focus on statistical reducibility: ${\mathbf{f}}$ is reducible if it is composed of two statistically independent co-occurring features (in which case ${\mathbf{f}}$ is “separable”) or if it is composed of two non-co-occurring features (in which case ${\mathbf{f}}$ is a “mixture”). We compare this with another way to define multi-dimensional irreducibility in Appendix C.
The probability distribution over input tokens $t$ induces a $d_{f}$-dimensional probability distribution over feature vectors ${\mathbf{f}}(t)$Fig. 2 shows two examples. Note that ${\mathbf{f}}(t)$ is a random vector since $t$ is a random variable; we use $p({\mathbf{f}})$ to denote the probability density function of ${\mathbf{f}}(t)$.
@@
    $p$ is a mixture $p({\mathbf{a}},{\mathbf{b}})=wp_{1}({\mathbf{a}},{\mathbf{b}})+(1-w)p_{2}({\mathbf{a}},{\mathbf{b}})$ of disjoint probability distributions for $w>0$, and $p_{1}$ is lower-dimensional such that $p_{1}({\mathbf{a}},{\mathbf{b}})=p_{1}({\mathbf{a}})\delta({\mathbf{b}})$.
Here $\delta$ is the Dirac delta function. By two probability distributions being disjoint, we mean that they have disjoint support (there is no set where both have positive probability measure, or equivalently the two features ${\mathbf{a}}$ and ${\mathbf{b}}$ cannot be active at the same time). In Eq. 1, ${\mathbf{a}}$ is the first $k$ components of the vector ${\mathbf{R}}{\mathbf{f}}+{\mathbf{c}}$ and ${\mathbf{b}}$ is the remaining $d_{f}-k$ components. When $p$ is separable or a mixture, we also say that ${\mathbf{f}}$ is separable or a mixture. We term a feature irreducible if it is not reducible, i.e., if no rotation and translation makes it separable or a mixture.
Fig. 2a shows an example of a 2D feature that is a mixture, because it can be decomposed into features ${\mathbf{a}}$ and ${\mathbf{b}}$ where ${\mathbf{b}}$ is a 1D line distribution (marked in red) and ${\mathbf{a}}$ is the remainder (a 2D cloud and a line, which can in turn be decomposed). An example of a feature that is separable is a normal distribution (since any multidimensional Gaussian can be rotated to have a diagonal covariance matrix). In natural language, a mixture might be a one hot encoding of “language of the current token”, while a separable distribution might be the “latitude” and “longitude” of location tokens.
In practice, because of noise and finite sample size, the mixture and separability definitions may not be precisely satisfied. Thus, we soften our definitions to permit degrees of reducibility:
@@
where $I$ denotes the mutual information. Smaller values of $S({\mathbf{f}})$ mean that ${\mathbf{f}}$ is more separable and therefore more reducible. Note that we can solely minimize over how many components to split off as ${\mathbf{a}}$ and over orthonormal matrices ${\mathbf{R}}$, since the additive offset ${\mathbf{c}}$ does not affect the mutual information.
@@
Larger values of $M_{\epsilon}({\mathbf{f}})$ mean that ${\mathbf{f}}$ is more of a mixture and is therefore more reducible.
We develop an optimization procedure to solve for the $\min$ and $\max$ in Definition 3 and apply it to synthetic feature distributions in Fig. 2a. Further details on this optimization process, more synthetic feature distributions, and more intuition about these definitions can be found in Appendix B.
@@ 3.2 Superposition
We now examine the implications of multi-dimensional features for the superposition hypothesis [11]. We first restate the original superposition hypothesis using our above definitions:
@@
Hypothesis 1 (One-Dimensional Superposition Hypothesis, paraphrased from [11]).
@@
We now present our new superposition hypothesis, which posits independence between irreducible multi-dimensional features instead of unknown levels of independence between one-dimensional features:
@@
The Johnson-Lindenstrauss (JL) Lemma [23] implies that we can choose $e^{Cd\delta^{2}}$ pairwise one-dimensional $\delta$-orthogonal vectors to satisfy 1 for some constant $C$, thus allowing us to build the model’s feature space with a number of one-dimensional $\delta$-orthogonal features exponential in $d$. In Appendix A, we prove a similar result for low-dimensional projections (the main idea of the proof is to combine $\delta$-orthogonal vectors as guaranteed from the JL lemma):
Theorem 1.
For any $d$, $d_{\max}$, and $\delta$, it is possible to choose $\frac{1}{d_{\max}}e^{C(d/d_{\max}^{2})\delta^{2}}$ pairwise $\delta$-orthogonal matrices ${\mathbf{A}}_{i}\in\mathbb{R}^{n_{i}\times d}$ where $1\leq n_{i}\leq d_{\max}$ for some constant $C$.
This exponential reduction in the number of features that are representable $\delta$-orthogonally (as opposed to a multiplicative factor reduction, which is also present) suggests that models will employ higher-dimensional features only for representations that necessitate detailed multi-dimensional descriptions. Moreover, these representations will be highly compressed to fit within the smallest dimensional space possible, potentially leading to interesting encodings; for example, recent work [33] finds that maximum-margin solutions for problems like modular arithmetic consist of Fourier features. Note that the proof assumes the “worst case” scenario that all of the features are dimension $d_{\max}$, while in practice many of the features may be $1$ or low dimensional, so the effect on the capacity of a real model that represents multi-dimensional features is unlikely to be this extreme.

### v2

@@ Abstract
Recent work has proposed that language models perform computation by manipulating one-dimensional representations of concepts (“features”) in activation space. In contrast, we explore whether some language model representations may be inherently multi-dimensional. We begin by developing a rigorous definition of irreducible multi-dimensional features based on whether they can be decomposed into either independent or non-co-occurring lower-dimensional features. Motivated by these definitions, we design a scalable method that uses sparse autoencoders to automatically find multi-dimensional features in GPT-2 and Mistral 7B. These auto-discovered features include strikingly interpretable examples, e.g. circular features representing days of the week and months of the year. We identify tasks where these exact circles are used to solve computational problems involving modular arithmetic in days of the week and months of the year. Next, we provide evidence that these circular features are indeed the fundamental unit of computation in these tasks with intervention experiments on Mistral 7B and Llama 3 8B. Finally, we find further circular representations by breaking down the hidden states for these tasks into interpretable components, and we examine the continuity of the days of the week feature in Mistral 7B.
@@ 1 Introduction
Language models trained for next-token prediction on large text corpora have demonstrated remarkable capabilities, including coding, reasoning, and in-context learning (Bubeck et al., 2023; Achiam et al., 2023; Anthropic, 2024; Team et al., 2023). However, the specific algorithms models learn to achieve these capabilities remain largely a mystery to researchers; we do not understand how language models write poetry. Mechanistic interpretability is a field that seeks to address this gap by reverse-engineering trained models from the ground up into variables (features) and the programs (circuits) that process these variables (Olah et al., 2020).
One mechanistic interpretability research direction has focused on understanding toy models in detail. This work has found multi-dimensional representations of inputs such as lattices (Michaud et al., 2024) and circles (Liu et al., 2022; Nanda et al., 2023a), and has successfully reverse-engineered the algorithms that models use to manipulate these representations. A separate direction has identified one-dimensional representations of high level concepts and quantities in large language models (Gurnee & Tegmark, 2023; Marks & Tegmark, 2023; Heinzerling & Inui, 2024; Bricken et al., 2023). These findings have led to the linear representation hypothesis (LRH): a hypothesis which has historically claimed both that 1. all representations in pretrained large language models lie along one-dimensional lines, and 2. model states are a simple sparse sum of these representations (Park et al., 2023; Bricken et al., 2023). In this work, we specifically call into question the first part of the LRH. [^1] An earlier version of this manuscript sparked discussion in the mechanistic interpretability community on the distinction between non-linear features and multi-dimensional features, and in fact this discussion directly led to a consensus around the two part LRH we describe above (see (Olah, 2024; Csordás et al., 2024; Mendel, 2024)). To clarify, we agree with these discussions, and believe the multi-dimensional features we find are ”linear” in the sense that they are contained in a low-dimensional linear subspace, but ”non-linear” in the sense that this low-dimensional subspace is not one-dimensional (and this is the sense we mean in the title).
For the most part, these two directions have been disconnected: Yedidia (2023a) and Gould et al. (2023) find intriguing hints of circular language model representations, and Bricken et al. (2023) speculate about the existence of feature manifolds, but these brief results only serve to further emphasize the lack of a unifying and satisfying perspective on the nature of language model features. In this work, we seek to bridge this gap by formalizing, investigating, and systematically searching for multi-dimensional language model features.
@@ 1.1 Contributions
    In Section 3, we generalize the one-dimensional definition of a language model feature to multi-dimensional features and provide an updated multi-dimensional superposition hypothesis to account for these new features.
@@
    In Section 4, we build on the definitions proposed in Section 3 to develop a theoretically grounded and empirically practical test that uses sparse autoencoders to find irreducible features. Using this test, we identify multi-dimensional representations automatically in GPT-2 and Mistral 7B, including circular representations for the day of the week and month of the year.
@@
    In Section 5, we show that Mistral 7B and Llama 3 8B use these circular representations when performing modular addition in days of the week and in months of the year. To the best of our knowledge, we are the first to find causal circular representations of concepts in a language model. We further investigate circular representations in these tasks by using regression to reveal circles in the computed day of the week and month of the year, and we additionally find that the model’s circular representations respect a continuous notion of time.
@@ 2 Related Work
Linear Representations: Early word embedding methods such as GloVe and Word2vec, although only trained using co-occurrence data, contained directions in their vector spaces corresponding to semantic concepts (Mikolov et al., 2013b; Pennington et al., 2014; Mikolov et al., 2013a). Recent research has found similar evidence of one-dimensional linear representations in sequence models trained only on next token prediction, including Othello board positions (Nanda et al., 2023b; Li et al., 2022), the truth value of assertions (Marks & Tegmark, 2023), and numeric quantities such as longitude, latitude, birth year, and death year (Gurnee & Tegmark, 2023; Heinzerling & Inui, 2024). These results have inspired the linear representation hypothesis (Park et al., 2023; Elhage et al., 2022) defined above. Jiang et al. (2024) provide theoretical evidence for this hypothesis, assuming a latent (binary) variable-based model of language. Empirically, Bricken et al. (2023) and Cunningham et al. (2023) successfully use sparse autoencoders to break down a model’s feature space into an over-complete basis of linear features. These works assume that the number of linear features stored in superposition exceeds the model dimensionality (Elhage et al., 2022).
Multi-Dimensional Representations: There has been comparatively little research on multi-dimensional features in language models. Yedidia (2023a; b) finds that GPT-2 learned position vectors form a helix, which implies a circle when “viewed” from below. Thus, we are not the first to find a circular feature in a language model. However, our work finds circular features that represent latent concepts from text, while the GPT-2 learned position vectors are specific to tokenization, separate from the rest of the model parameters, and causally implicated only due to positional attention masking. Another suggestive result, due to Hanna et al. (2024), is the presence of a U-shape in the representation of numbers between 00 and $100$; however, Hanna et al. (2024) find that this representation is not causal, and they only show it exists within a specific prompt distribution. Recent work on dictionary learning (Bricken et al., 2023) has speculated about multi-dimensional feature manifolds; our work is similar to this direction and develops the idea of feature manifolds theoretically and empirically. Finally, in a separate direction, Black et al. (2022) argue for interpreting neural networks through the polytopes they split the input space into, and identifies regions of low polytope density as “valid” regions for a potential linear representation.
@@
Circuits: Circuits research seeks to identify and understand circuits, subsets of a model (usually represented as a directed acyclic graph) that explain specific behaviors (Olah et al., 2020). The base units that form a circuit can be layers, neurons (Olah et al., 2020), or sparse autoencoder features (Marks et al., 2024). In the first circuits-style work, Olah et al. (2020) found line features that were combined into curve detection features in the InceptionV1 image model. More recent work has examined language models, for example the indirect object identification circuit in GPT-2 (Wang et al., 2022). Given the difficulty of designing bespoke experiments, there has been increased research in automated circuit discovery methods (Marks et al., 2024; Conmy et al., 2023; Syed et al., 2023).
@@
Interpretability for Arithmetic Problems: Liu et al. (2022) study models trained on modular arithmetic problems $a+b=c\ (\mathrm{mod}\ m)$ and find that models that generalize well have circular representations for $a$ and $b$. Further work by Nanda et al. (2023a) and Zhong et al. (2024) shows that models use these circular representations to compute $c$ via a “clock” algorithm and a separate “pizza” algorithm. These papers are limited to the case of a small model trained only on modular arithmetic. Another direction has studied how large language models perform basic arithmetic, including a circuits level description of the greater-than operation in GPT-2 (Hanna et al., 2024) and addition in GPT-J (Stolfo et al., 2023). These works find that to perform a computation, models copy pertinent information to the token before the computed result and perform the computation in the subsequent MLP layers. Finally, recent work by Gould et al. (2023) investigates language models’ ability to increment numbers and finds linear features that fire on tokens equivalent modulo $10$.
3 Definitions
@@
This section focuses on hypotheses for how hidden states of language models can be decomposed into sums of functions of the input (features). We focus on $L$ layer transformer models $M$ that take in token input ${\bf t}=(t_{1},\ldots,t_{n})$, have hidden states $\mathbf{x}_{1,l},\ldots,\mathbf{x}_{n,l}$ for layers $l$, and output logit vectors $\mathbf{y}_{1},\ldots,\mathbf{y}_{n}$. Given a set of inputs $T$, we let $X_{i,l}$ be the set of all corresponding $\mathbf{x}_{i,l}$. We write matrices in capital bold, vectors and vector valued functions in lowercase bold, and sets in capital non-bold.
@@ 3.1 Multi-Dimensional Features
We define a $d_{f}$-dimensional feature as a function ${\mathbf{f}}$ that maps a subset of the input space into a $d_{f}$-dimensional point cloud in $\mathbb{R}^{d_{f}}$. We say that a feature is active on the aforementioned subset.
The probability distribution over input tokens $t$ induces a $d_{f}$-dimensional probability distribution over feature vectors ${\mathbf{f}}(t)$. As an example, let $n=1$ (so inputs are single tokens) and consider a feature ${\mathbf{f}}$ that maps integer tokens to their integer values in $\mathbb{R}^{1}$. Then ${\mathbf{f}}$ is a $1$-dimensional feature that is active on integer tokens, and ${\mathbf{f}}(t)$ is the marginal integer occurrence distribution from the token distribution.
How can we differentiate "true" multi-dimensional features from sums of lower dimensional features? We make this distinction by examining the reducibility of a potential multi-dimensional feature. That is, ${\mathbf{f}}$ is a "true" multi-dimensional feature if it cannot be written as the sum of two statistically independent features and it cannot be written as the sum of two non-co-occurring features. Formally, we have the following definition:
@@
    $p$ is a mixture, i.e., a sum of disjoint distributions, one of which is lower dimensional:
    $p({\mathbf{a}},{\mathbf{b}})=wp({\mathbf{a}})\delta(b)+(1-w)p({\mathbf{a}},{\mathbf{b}})$
Here $\delta$ is the Dirac delta function and $10<w<$ 1. By two probability distributions being disjoint, we mean that they have disjoint support (there is no set where both have positive probability measure, or equivalently the two features ${\mathbf{a}}$ and ${\mathbf{b}}$ cannot be active at the same time). In Eq. 1, ${\mathbf{a}}$ is the first $k$ components of the vector ${\mathbf{R}}{\mathbf{f}}+{\mathbf{c}}$ and ${\mathbf{b}}$ is the remaining $d_{f}-k$ components. When $p$ is separable or a mixture, we also say that ${\mathbf{f}}$ is separable or a mixture. We term a feature irreducible if it is not reducible, i.e., if no rotation and translation makes it separable or a mixture.
An example of a feature that is a mixture is a one hot encoding along a simplex; an example of a feature that is separable is a normal distribution222since any multidimensional Gaussian can be rotated to have a diagonal covariance matrix. In natural language, a mixture might be a one hot encoding of “breed of dog”, while a separable distribution might be the “latitude” and “longitude” of location tokens.
In practice, the mixture and separability definitions may not be precisely satisfied. Thus, we soften our definitions to permit degrees of reducibility:
@@
where $I$ denotes the mutual information. Smaller values of $S({\mathbf{f}})$ mean that ${\mathbf{f}}$ is more separable.
@@
Larger values of $M_{\epsilon}({\mathbf{f}})$ mean that ${\mathbf{f}}$ is more of a mixture.
In Appendix B, we expand on the intuition behind why the separability and $\epsilon$-mixture indices correspond to weakened versions of Definition 2. In Section B.2, we also describe an optimization procedure we developed to solve for the $\min$ and $\max$ in Definition 3 in practice. In Section 4, we apply this empirical test to real language model feature distributions to find irreducible multi-dimensional features.
@@ 3.2 Superposition
In this section, we propose an updated superposition hypothesis (Elhage et al., 2022) that takes into account multi-dimensional features. First, we restate the original superposition hypothesis:
@@
Hypothesis 1 (One-Dimensional Superposition Hypothesis, paraphrased from (Elhage et al., 2022)).
@@
In contrast, our new superposition hypothesis posits independence between irreducible multi-dimensional features instead of unknown levels of independence between one-dimensional features:
@@
Note that since multi-dimensional features can be written as the sums of projections of lower-dimensional features, our new superposition hypothesis is a stricter version of 1. In the next section, we will explore empirical evidence for our hypothesis, while in Appendix A, we prove upper and lower bounds on the number of $\delta$-almost orthogonal matrices $\mathbf{V}_{i}$ that can be packed into $d$ dimensional space.
Figure 2: Empirical $\epsilon$-mixture index and separability index for the “days of the week” cluster along PCA components 2 and 3. This point cloud has a lower $\epsilon$-mixture index and higher separability index than PCA projections within typical clusters, indicating that it is more likely to be an irreducible multi-dimensional feature.